JÁNOS BOLYAI MATHEMATICAL SOCIETY Budapest, Fő utca 68., 1027

phone: 201 - 7656, phone/fax: 201 - 6974, e-mail: H3341sza@ella.hu

INTERNATIONAL COLLOQUIUM ON COMBINATORICS AND GRAPH THEORY A satellite conference of ECM2 Balatonlelle, Hungary, July 15-20, 1996

Dear Colleagues,

The J. Bolyai Mathematical Society organizes a combinatorial conference entitled ”International Colloquium on Combinatorics and Graph Theory” at Balatonlelle, Hungary, on July 15-20, 1996 the week before ECM2. The purpose of this conference is to discuss recent advances in Combinatorics. The following 9 sections are planned:

- computational geometry

- combinatorial optimization

- combinatorial set theory

- combinatorial number theory

- graph theory

- extremal set theory & symmetric structures

- theoretical computer science

- probabilistic methods

- combinatorial methods of molecular biology.

The conference will be held in Balatonlelle, a resort area about 140 km (90 miles) from Budapest, in a hotel with a private beach. A preliminary information about the prices (which are subject to change, unfortunately):

- accommodation – appr. 16 - 20 USD/night/person (with breakfast),

- lunch or dinner – appr. 5 - 8 USD, each. (Railway fare: appr. 13 USD; round trip ticket.) Enclosed please find the registration form of the conference, which should be filled in and returned to our address above by February 29, 1996. E-mail registration is strongly suggested. Please visit our Web page at

    http://www.math-inst.hu/conferences/comb96/home.html,

where the registration form is placed. The conference’s e-mail address is: comb96@math-inst.hu We will send you more detailed information in late March.

We are looking forward to seeing you in Hungary.

Sincerely yours,

L. Lovász (chair) Z. Füredi, E. Győri (co-chairs)

M. Ruszinkó, A. Sali (secretaries),

I. Bárány, J. Beck, B. Bollobás, A. Frank, A. Gyárfás, A. Hajnal, G.O.H. Katona, P. Komjáth, J. Komlós, J. Pach, A. Recski, M. Simonovits, E. Szemerédi, L. Székely, É. Tardos, Zs. Tuza, members of the O.C.

January 8, 1996 Enclosures